### v1

# The Coverage Principle: A Framework for Understanding Compositional Generalization

Abstract
Large language models excel at pattern matching, yet often fall short in systematic compositional generalization. We propose the coverage principle: a data-centric framework showing that models relying primarily on pattern matching for compositional tasks cannot reliably generalize beyond substituting fragments that yield identical results when used in the same contexts. We demonstrate that this framework has a strong predictive power for the generalization capabilities of Transformers. First, we derive and empirically confirm that the training data required for two-hop generalization grows at least quadratically with the token set size, and the training data efficiency does not improve with 20x parameter scaling. Second, for compositional tasks with path ambiguity where one variable affects the output through multiple computational paths, we show that Transformers learn context-dependent state representations that undermine both performance and interpretability. Third, Chain-of-Thought supervision improves training data efficiency for multi-hop tasks but still struggles with path ambiguity. Finally, we outline a mechanism-based taxonomy that distinguishes three ways neural networks can generalize: structure-based (bounded by coverage), property-based (leveraging algebraic invariances), and shared-operator (through function reuse). This conceptual lens contextualizes our results and highlights where new architectural ideas are needed to achieve systematic compositionally. Overall, the coverage principle provides a unified lens for understanding compositional reasoning, and underscores the need for fundamental architectural or training innovations to achieve truly systematic compositionality. [^1] All code and resources are available at: https://github.com/kaistAI/coverage-principle
1 Introduction
[…]
Figure 1: Left: In a two-hop task $(x_{1},x_{2},x_{3})\mapsto t$ with $t=f_{2}(f_{1}(x_{1},x_{2}),x_{3})$, two fragments $(x_{1},x_{2})$ and $(x_{1}^{\prime},x_{2}^{\prime})$ satisfying $f_{1}(x_{1},x_{2})=f_{1}(x_{1}^{\prime},x_{2}^{\prime})=b$ consistently yield the same final output when combined with the same context $x_{3}$, supporting their functional equivalence. Right: Among all possible inputs (few shown), we draw an edge between any two inputs that differ only by functionally equivalent fragments to form a substitution graph. Then, coverage is the set of observed inputs (highlighted as blue) and all inputs connected to them. The coverage principle states that a learner relying on pattern matching can reliably generalize only to inputs that are reachable from the observations by substituting functionally equivalent fragments.
Human generalization capability is often attributed to our ability to manipulate symbols compositionally [1, 2]. Despite the remarkable performance of Large Language Models (LLMs) [3, 4, 5, 6, 7, 8], our understanding of how they represent and manipulate compositional structures is still limited. Meanwhile, a growing body of work points to the brittleness of LLMs in performing systematic [^2] We refer to systematicity as an ability to generalize to unseen outcomes by exploiting their structural relations. For a more thorough discussion, see Fodor and Pylyshyn [9], Marcus [10], and Szabó [11]. compositional reasoning [12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25]. These observations suggest that LLMs succeed primarily through pattern matching, i.e., by exploiting surface-level statistical correlations between input fragments and outputs, rather than systematic compositional reasoning. However, prior analyses of the pattern-matching behavior remain largely observational and task-specific, leaving us without a unified framework to predict when they will succeed or fail.
We show that Transformers’ [26] generalization on synthetic compositional tasks is predictable through a unified framework grounded in one observation: Transformers generalize by utilizing functionally equivalent fragments supported by input-output relations. This insight allows us to move beyond observing success or failure cases to systematically predicting generalization boundaries. To this end, we first formalize the concept of functional equivalence, where two fragments (i.e., components of input sequences) are regarded as equivalent when they yield identical results in all shared contexts observed in the training data. Then, we introduce coverage, the set of inputs reachable from training data by substituting functionally equivalent fragments. Finally, we propose the coverage principle, which states that coverage is a necessary condition for compositional generalization, i.e., model predictions become unreliable beyond this fundamental boundary, as long as the only generalization mechanism is pattern matching.
From this conceptual formulation, we derive specific predictions about generalization in Transformers. First, we derive and empirically confirm that reliable generalization on a two-hop reasoning task requires training data to scale at least quadratically with the token set size. Remarkably, our experiments suggest that this scaling relationship remains invariant even for 20x larger model, underscoring that the limitation is inherent to data properties rather than model capacity.
Second, our framework also predicts specific complications in tasks with path ambiguities, where the same variable influences outputs through multiple computational paths. We show that Transformers cannot form unified representations of identical intermediate states in such cases, instead developing context-dependent state representations that fragment across different paths. This degrades both performance and interpretability, explaining why seemingly simple compositional tasks can remain challenging even for large models.
Third, while Chain-of-Thought (CoT) supervision [27, 28] improves data efficiency by ‘flattening’ tasks into sequences of single-hop problems, our analysis reveals a fundamental limitation: CoT-trained models still cannot resolve path ambiguities without showing nearly all possible in-domain combinations. This finding provides insights into understanding the Transformers’ failure modes in complex reasoning tasks with path ambiguities (e.g., planning [29, 30, 31]).
We place these findings within a broader, mechanism-based taxonomy that distinguishes three routes to generalization: (i) structure-based generalization via functional equivalence (the focus of this paper), (ii) property-based generalization that exploits algebraic invariances of individual primitive functions, and (iii) shared-operator generalization that reuses the same computation across positions. This taxonomy clarifies how certain model behaviours that occur outside the coverage region can arise from Type-II or Type-III mechanisms, providing a conceptual framework for analysing diverse sources of generalization.
Overall, the coverage principle offers a unified conceptual framework providing valuable insights into compositional generalization, highlighting how compositional reasoning and its mechanistic interpretation remain inherently bounded by coverage limitations. This underscores the need for architectural or training innovations to achieve truly systematic generalization in neural networks.
2 Related work
Compositional generalization and systematicity
Since Fodor and Pylyshyn [9] posed the systematicity challenge, arguing connectionist systems inherently lack symbolic structural capabilities, understanding how neural networks might achieve compositional reasoning has remained a persistent research question [32, 33, 34, 35, 36, 37, 38, 39, 10]. Central to this debate is the variable binding problem, how neural networks can flexibly associate variables with values across contexts [40, 41, 42], which Marcus [37] argues as the fundamental limitation of modern connectionist approaches. Despite AI advances, a growing body of work reveals limitations in neural networks’ compositional capabilities [14, 43, 18, 44, 45, 20, 46, 15, 21, 22, 23, 24, 47, 42]. Dziri et al. [17] observe that Transformers perform linearized subgraph matching to solve compositional tasks, but leave open when and why it occurs. Our coverage principle addresses this gap by formally demarcating what pattern matching models can achieve without true variable binding, showing both the possibilities and fundamental limitations of pattern matching-based approaches through data-based preconditions. This aligns with empirical findings of Andreas [48] and Auersperger and Pecina [49], where compositional data augmentation that implicitly expands coverage significantly improves generalization. Our derived power-law scaling relationship between token set size and required dataset size contributes to scaling laws literature [50, 51], but differs by characterizing a data requirement invariant to model size for compositional tasks. Unlike prior disentanglement work [52, 53, 54, 55] which focuss on latent factors, our data-centric view identifies fundamental boundary conditions for pattern-based generalization. The coverage principle thus complements recent architectural or training innovations [56, 57, 58, 59, 60] by showing what any pattern matching approach must overcome to achieve true systematicity.
Mechanistic interpretability
Mechanistic interpretability studies aim to understand how submechanisms implement models’ behaviors [61, 62, 63, 64]. Recent work traces how Transformer components coordinate on compositional tasks [65, 66, 67, 68, 69]. Wang et al. [19] demonstrate that in-domain compositional generalization emerges through grokking, but do not identify necessary data conditions for such emergence. Our data-centric framework complements this optimization-centric view by pinpointing when generalizing circuits can exist: only when training data provides sufficient evidence of functional equivalence, essentially formalizing when the pattern matching mechanisms that Fodor and Pylyshyn [9] critique can succeed or must fail. Our findings also explain why standard interpretability techniques like logit lens [70, 71] may fail to identify compositional structure in models trained on tasks with path ambiguities.
3 A formulation of the coverage principle
We now develop a formal framework of the coverage principle. We first provide an intuitive illustration with a two-hop compositional task, then generalize the intuitions to formalize the coverage principle for more general compositional tasks.
Imagine a learner [^3] By “learner”, we mean any system that learns from data (e.g., neural networks). observing input-output pairs determined by a function $f:{\mathcal{X}}^{3}\to{\mathcal{X}}$. The input ${\bm{x}}=(x_{1},x_{2},x_{3})\in{\mathcal{X}}^{3}$ is a sequence of three discrete tokens and the output is a single token, where each token is sampled from a finite set ${\mathcal{X}}$ [^4] For brevity, we use a shared token set ${\mathcal{X}}$. Position-specific domains ${\mathcal{X}}_{i}$ can be embedded as subsets of an enlarged token set $\tilde{{\mathcal{X}}}={\mathcal{X}}_{1}\cup{\mathcal{X}}_{2}\cup{\mathcal{X}}_{3}$ without loss of generality. Being unknown to the learner, let $f$ be a function that can be factorized as the composition of two primitive functions as $f({\bm{x}})=f_{2}(f_{1}(x_{1},x_{2}),x_{3})$, where $f_{1}:{\mathcal{X}}^{2}\to{\mathcal{X}}$ and $f_{2}:{\mathcal{X}}^{2}\to{\mathcal{X}}$, as illustrated in Fig. 2(a). How can the learner generalize by only seeing the input-output pairs?
Our key intuition is that a learner exploits the compositional structure only when observations show that two fragments behave identically. For instance, assume that $f_{1}(x_{1},x_{2})=f_{1}(x_{1}^{\prime},x_{2}^{\prime})=b$ for $(x_{1},x_{2}),(x_{1}^{\prime},x_{2}^{\prime})\in{\mathcal{X}}^{2}$. From the underlying structure, these fragments are functionally equivalent, i.e., substituting one for the other never changes the output. If observations consistently support their equivalence, i.e., $f(x_{1},x_{2},x_{3})=f(x_{1}^{\prime},x_{2}^{\prime},x_{3})$ for observed $x_{3}$ values, this equivalence can be supported (Fig. 1 (Left)). Intuitively, the learner would be able to harness this equivalence to generalize to an unseen input $(x_{1}^{\prime},x_{2}^{\prime},x_{3}^{\prime\prime})$, provided the learner has seen $f(x_{1},x_{2},x_{3}^{\prime\prime})$.
In short, the learner can utilize the observed functional equivalence to correctly infer the output of an unseen input, if it can reach an observed input by safe substitutions supported by observations (Fig. 1 (Right)). Coverage is a set of such inputs that are reachable from an observed input through chains of functionally equivalent substitutions. The essence of our framework, the coverage principle, is that a learner can only generalize inside the coverage when the only source of compositional generalization is the observation and utilization of functional equivalence.
We now formalize these concepts for an arbitrary fixed-length task with an arbitrary set of observations. We restrict our attention to single-token prediction tasks defined as a deterministic mapping $f:{\mathcal{X}}^{n}\to{\mathcal{X}}$, where ${\mathcal{X}}$ is a finite set of tokens. We also consider a fixed observation set $D\subset{\mathcal{X}}^{n}$, a collection of inputs that are allowed to be observed by the learner. Write ${\bm{x}}=(x_{1},\dots,x_{n})\in{\mathcal{X}}^{n}$ and, for a subset $I\subset[n]:=\{1,\dots,n\}$, let ${\bm{x}}_{I}:=(x_{i})_{i\in I}$ be a subsequence of ${\bm{x}}$. The first step is to formalize what it means for two subsequences to be functionally equivalent.
Definition 3.1 (Functional $k$-equivalence).
Fix a nonempty proper subset of indices $I\subsetneq[n]$. Given a pair of subsequences ${\bm{a}},{\bm{a}}^{\prime}\in{\mathcal{X}}^{|I|}$, we say a pair of inputs $\left\{{\bm{x}},{\bm{x}}^{\prime}\right\}\subset D$ to be an $I$-co-occurrence of ${\bm{a}}$ and ${\bm{a}}^{\prime}$ in $D$ if it satisfies ${\bm{x}}_{I}={\bm{a}}$, ${\bm{x}}^{\prime}_{I}={\bm{a}}^{\prime}$, and ${\bm{x}}_{[n]\setminus I}={\bm{x}}^{\prime}_{[n]\setminus I}$. Also, we say ${\bm{a}}$ and ${\bm{a}}^{\prime}$ are functionally $k$-equivalent at $I$ in $D$, denoted by ${\bm{a}}\equiv^{I}_{D}{\bm{a}}^{\prime}$, if it satisfies:
  1. 1.
    (Sufficiency of co-occurrences.) There are $k$ or more distinct $I$-co-occurrences of ${\bm{a}}$ and ${\bm{a}}^{\prime}$ in $D$;
  2. 2.
    (Consistency.) Every $I$-co-occurrence $\left\{{\bm{x}},{\bm{x}}^{\prime}\right\}$ of ${\bm{a}}$ and ${\bm{a}}^{\prime}$ in $D$ satisfies $f({\bm{x}})=f({\bm{x}}^{\prime})$.
In other words, two subsequences are functionally $k$-equivalent if they behave identically in the same contexts at least $k$ times. The hyperparameter $k$ represents the strength of evidence required to establish functional equivalence between two subsequences. The minimum value $k=1$ corresponds to the weakest form of evidence, meaning a single shared context is sufficient to establish equivalence, whereas higher values of $k$ demand more robust evidence.
Next, we ask: which inputs are reachable from observed data utilizing functional equivalence? To formalize this, we define the substitution graph: Let ${\mathcal{G}}_{D,k}=(V,E)$ be an undirected graph with a vertex set $V={\mathcal{X}}^{n}$ of all possible inputs. Two vertices ${\bm{x}},{\bm{x}}^{\prime}\in V$ are connected with an edge in $E$ if and only if there exists an index set $I\subset[n]$ such that $\left\{{\bm{x}},{\bm{x}}^{\prime}\right\}$ is an $I$-co-occurrence (in $V$) of a pair of functionally $k$-equivalent sequences at $I$ in $D$. With this substitution graph ${\mathcal{G}}_{D,k}$, we formally define the $k$-coverage as a set of inputs which are connected [^5] For an undirected graph ${\mathcal{G}}$, two vertices $u,v$ are connected if ${\mathcal{G}}$ contains a path between $u$ and $v$. to at least one observed input as follows:
Definition 3.2 ($k$-coverage).
The $k$-coverage of $D$, denoted by $\mathrm{Cover}_{k}(D)$, is the set of all inputs in ${\mathcal{X}}^{n}$ that is connected to an input in $D$ on the substitution graph ${\mathcal{G}}_{D,k}$.
Note that the notion of coverage is a stricter condition of the conventional definition of in-domain, which is obtained by random train/test split [19] or taking combinations of observed internal computations [17]. We also emphasize that coverage is a property of a dataset and is independent of model architectures and learning algorithms. Now, we formally phrase the coverage principle:
Claim 3.3 (Coverage principle).
Let a learner output $\hat{f}$ that is consistent with the training set $D$. If generalization is derived solely from observed functional $k$-equivalences and ${\bm{x}}\notin\mathrm{Cover}_{k}(D)$, the value $\hat{f}({\bm{x}})$ is unconstrained by $D$.
This establishes a fundamental boundary: no learner can reliably generalize beyond the coverage as long as the only source of the generalization is functional equivalence. In Sec. 5.1, we show that Transformers indeed fail to generalize beyond coverage when our random mapping dataset construction isolates functional equivalence as the sole generalization source. However, deeper insights come from asking when generalization success occurs in practice. While theory suggests that deep learning methods implicitly favor simpler solutions [72, 73, 74], which may involve leveraging functional equivalence, we still lack a systematic picture of how task structure, dataset, and model size interact to determine when such solutions emerge. We study this through experiments in Sec. 5.
4 Experimental setup
$x_{1}$$x_{2}$$f_{1}$$b$$x_{3}$$f_{2}$$t$
(a) 2-Hop
$x_{1}$$x_{2}$$x_{3}$$x_{4}$$f_{1}$$f_{2}$$b_{1}$$b_{2}$$f_{3}$$t$
(b) Parallel 2-Hop
$x_{1}$$x_{2}$$f_{1}$$b_{1}$$x_{3}$$f_{2}$$b_{2}$$x_{4}$$f_{3}$$t$
(c) 3-Hop
$x_{1}$$x_{2}$$f_{1}$$b$$x_{3}$$f_{2}$$t$
(d) Non-Tree
Figure 2: Four synthetic compositional structures we study.
Dataset construction
We construct four synthetic compositional tasks: 2-Hop, Parallel 2-Hop, 3-Hop, and Non-tree (Fig. 2). We create random mappings from the product space of a token set to control generalization sources not attributable to compositional structures (i.e., commutativity). [^6] Such external generalization sources will be discussed in Sec. 7 and App. H. We explain the dataset construction process using 2-Hop task (Fig. 2(a)), $(x_{1},x_{2},x_{3})\mapsto t$ with $t=f_{2}(f_{1}(x_{1},x_{2}),x_{3})$, as an example. We construct training datasets by defining a token set with size $|{\mathcal{X}}|$, and creating two random maps for the primitive functions $f_{1}:{\mathcal{X}}^{2}\to{\mathcal{X}}$ and $f_{2}:{\mathcal{X}}^{2}\to{\mathcal{X}}$. We mark a fraction $p_{\mathrm{seen}}=0.7$ of each function’s domain as ‘seen’, gather all possible combinations where both functions are applied to inputs from their seen domains, and uniformly sample $N$ examples to form a training dataset. See Sec. B.1 for more details of the dataset construction process.
Training & evaluation
Following Wang et al. [19], we train randomly initialized GPT-2 [75] models with 8 layers, 12 heads, and 768 dimensions (see Sec. B.2 for details). We construct two evaluation sets, each with 2,000 instances: (1) In-Domain (ID) Test Set : all primitive function applications (e.g., $f_{1}(x_{1},x_{2})$ and $f_{2}(b,x_{2})$ in 2-Hop task) are observed during training, but the specific combination was unseen. (2) Out-of-Domain (OOD) Test Set: at least one primitive function application is never observed during training.
5 Results
5.1 Predictive power of $k$-coverage on Transformer’s compositional generalization
Figure 3: Left: Percentage of covered ID data depending on $k$ values and dataset size ($N$), for 2-Hop task ($|{\mathcal{X}}|=50$). Right: Test accuracy depending on $k$-cutoff values for 2-Hop task ($|{\mathcal{X}}|$=50, $N$=10k). Each line represents a different training checkpoint. Note that OOD accuracy remains at chance level (=1/50) regardless of training time. The bars below show the number of test data for each $k$-cutoff value.
We now empirically validate the predictive power of the coverage principle by quantifying the correlation between $k$-coverage and model generalization performance. To this end, we implement and release a task-agnostic coverage determination algorithm (see App. C for implementation details) that can be applied to diverse compositional structures. Then, we analyze what fraction of ID test data of 2-Hop task with $|{\mathcal{X}}|=50$ lies inside $k$-coverage, depending on $k$ and dataset size $N$. Fig. 3 (Left) shows that at $N=5$k, every ID test example is already covered with minimal evidence ($k=1$).
Ideally, a single witness of functional equivalence could suffice. However, experiments show that such minimal coverage alone is insufficient for efficient generalization in practice. To demonstrate this, we define a sample’s $k$-cutoff as the lowest $k$ for which an input lies in $k$-coverage, measuring the strength of evidence for functional equivalence. For example, a $k$-cutoff of 3 means that an example is inside coverage with $k=3$ but not with $k=4$. For 2-Hop dataset with $N=10$k, we classify each ID test instance according to its $k$-cutoff, and track accuracy development for each group across 50k training epochs. As shown in Fig. 3 (Right), generalization success shows a strong relationship with $k$-cutoff values. Test data with low $k$-cutoff values show delayed improvement even after extensive training, while examples with stronger evidence generalize much faster. Notably, OOD test examples (outside coverage) remain at chance-level ($1/50=0.02$) throughout training, empirically validating our coverage principle.
These results yield two important insights. First, successful generalization in practice requires robust coverage so the model can confidently identify and utilize functional equivalence relationships. The $k$ parameter effectively quantifies this evidence strength, directly impacting generalization speed and reliability. Second, while our experiments use uniformly distributed data, the coverage principle explains why models struggle with generalizing long-tail distribution in imbalanced real-world data [76, 77, 78]. Rare combinations naturally receive limited functional equivalence evidence (low $k$), placing them effectively outside practical coverage, despite technically being in-distribution.
5.2 The clustering of latent representation drives generalization on coverage
Figure 4: Left: Heatmap of Intra-Inter Cosine Gap (IICG) across layers and positions, sliced by $k$-cutoff. Higher IICG values indicate stronger clustering of representations that share the same intermediate state. The positions with the highest IICG values are marked with squares. Right: PCA visualization of latent representations at position $x_{2}$ and layer 3. Datapoints are classified by their intermediate states $b=f_{1}(x_{1},x_{2})$.
Next, we investigate how the model internally represents functional equivalence for $k$-covered inputs. Specifically, we inspect a GPT-2 trained on 2-Hop task ($|{\mathcal{X}}|=50$, $N=10\text{k}$) for 50k epochs (corresponding to the yellow line in Fig. 3 (Right)). [^7] Analyses for varying factors including task structures, entity set size ($|{\mathcal{X}}|$), dataset size ($N$), and training steps give consistent results; see App. D. We observe that when a model successfully generalizes to ID test data, it maps functionally equivalent components into tight latent clusters, thereby encoding the equivalence relationships needed for compositional generalization.
To quantify this representation clustering phenomenon, we develop a metric that captures how distinctly the model separates functionally equivalent fragments from others. Specifically, we measure the difference between the average pairwise cosine similarity of latent vectors that share the same intermediate state $b=f_{1}(x_{1},x_{2})$ ($\overline{\cos}_{\text{intra}}$), and those that do not ($\overline{\cos}_{\text{inter}}$), for each position and layer of the model. We term this difference the Intra–Inter Cosine Gap $\text{IICG}=\overline{\cos}_{\text{intra}}-\overline{\cos}_{\text{inter}}$, where higher values indicate stronger within-group clustering relative to between-group separation. Fig. 4 (Left) reveals a clear relationship: higher $k$-cutoff values yield higher IICG scores at certain positions, indicating that stronger functional equivalence evidence leads to more coherent internal representations. Critically, OOD examples show no clustering pattern at all, since they lack any functional equivalence evidence in the training data, as predicted by the coverage principle. The PCA visualization at position $x_{2}$ and layer 3 (Right) shows this trend visually. We verify that the representation clusters play a causal role in ID generalization with causal tracing [67, 66], a widely used mechanistic interpretability technique to identify Transformer circuits (see Fig. 8 in App. D).
Our findings extend the insights of Wang et al. [19] in several ways. First, we demonstrate that unified circuit formation is driven by functional equivalence evidence in the training data, not by explicit exposure to intermediate computation steps. Consequently, out-of-coverage inputs lack these unified circuits by construction, explaining their systematic failure on OOD examples. Moreover, we find that these clustered representations are not necessarily aligned with vocabulary embeddings, implying that standard interpretability methods like logit lens [70] may fail to detect these functional equivalence representations despite their presence. See App. I for a detailed analysis.
5.3 When does coverage lead to full ID generalization? A power-law lower bound
Figure 5: Left: Log-log plot of measured $\hat{N}_{\mathrm{req}}$ vs. token set size ($|{\mathcal{X}}|$) across three compositional tasks. The slope $c$ corresponds to the empirical power-law scaling exponent. Omitted points for 3-Hop are due to prohibitively large dataset requirements. Right: Power-law scaling remains invariant across GPT-2 model sizes (68M to 1.5B parameters) for 2-Hop task. $R^{2}>0.99$ for all linear fitting.
Our analysis of $k$-coverage and representation clustering demonstrates that stronger functional equivalence evidence leads to better generalization. A natural follow-up question arises: How large should the training set be to enable full generalization on all ID test data? Intuitively, this requires the training set to support the functional equivalence of every pair of inputs that shares the same intermediate state $b$. Formally, for a 2-Hop task we need $(x_{1},x_{2})\equiv^{\{1,2\}}_{D}(x^{\prime}_{1},x^{\prime}_{2})$ whenever $f_{1}(x_{1},x_{2})=f_{1}(x^{\prime}_{1},x^{\prime}_{2})$. Assuming a learner requires at least $k$ distinct pairs of evidence to establish functional equivalence of two fragments, how does the required dataset size scale with the token set size $|{\mathcal{X}}|$? In practical terms, this question addresses how much data is fundamentally needed for compositional generalization, which is central to understanding data scaling requirements for compositional tasks. For 2-Hop task, we derive the following result (full derivation in App. E):
Result 5.1 (Power-law lower bound).
Let $f:{\mathcal{X}}^{3}\!\to\!{\mathcal{X}}$ be a 2-Hop composition. Assume a learner recognizes functional equivalence of two subsequences only after observing them in at least $k$ distinct pairs of evidence (i.e., functionally $k$-equivalent). [^8] Note that this assumption is equivalent to modeling Fig. 3 (Right) as a step function. Let $N_{\text{req}}(|{\mathcal{X}}|,k)$ be the smallest training dataset size that enables complete generalization to ID examples under this evidence threshold. Then, up to poly-logarithmic factors in $|{\mathcal{X}}|$, $N_{\text{req}}(|{\mathcal{X}}|,k)=\tilde{\Omega}\bigl{(}|{\mathcal{X}}|^{\,\alpha(k)}\bigr{)},\text{where}\;\alpha(k)=2.5-\frac{0.5}{k}.$
Res. 5.1 predicts that a learner relying on pattern matching requires training dataset size scaling at least quadratically with respect to the token set size, to fully generalize on ID test data. To empirically confirm this, we define a practical threshold $\hat{N}_{\mathrm{req}}$ to estimate $N_{\mathrm{req}}(|{\mathcal{X}}|,k)$, as a minimal amount of training data required to exceed ID accuracy of 0.99 within 100 epochs after reaching the same level on training data. [^9] See App. F measurement details. Fig. 5 (Left) shows the measured power-law exponents for $\hat{N}_{\text{req}}$ vs. $|{\mathcal{X}}|$ across different task structures. The measured exponent for 2-Hop ($c=2.26$) aligns well with our theoretical predictions of at least quadratic scaling. Although we derive the theoretical bound only for 2-Hop, we observe clear power-law relationships for more complex structures as well. The higher exponents for Parallel-2-Hop ($c=2.43$) and 3-Hop ($c=2.58$) tasks suggest that extra computational steps essentially add another dimension of relationships that require robust coverage, driving the steeper power-law scaling.
Strikingly, these exponents remain invariant across three different GPT-2 model sizes spanning a 20x range in parameters (from 68M to 1.5B) for all three tasks (Fig. 5 (Right) and Tab. 2 in App. F). This supports that the scaling relationship is primarily determined by data properties rather than model capacity, reinforcing our framework’s data-centric perspective. The observed scaling relationships are robust across different hyperparameters (weight decay and learning rate) and empirical decision criteria for $\hat{N}_{\mathrm{req}}$ (see App. F). Overall, these results suggest that compositional generalization in Transformers requires dataset sizes scaling as a power-law with token set size. This has significant implications for how we approach complex reasoning tasks, suggesting that data curation strategies informed by coverage considerations may be more effective than raw parameter scaling.
5.4 Path ambiguity hinders generalization and interpretability in Transformers
Identical Intermediate StatesFunctional Equivalence$(x_{1},x_{2})$$(x_{1}^{\prime},x_{2})$$(x_{1},x_{2}^{\prime})$$(x_{1}^{\prime},x_{2}^{\prime})$$(x_{1},x_{2}^{\prime\prime})$$(x_{1}^{\prime},x_{2}^{\prime\prime})$
Figure 6: Left: In Non-tree task, the coverage principle predicts that representations of input subsequences with the same intermediate state $b=f_{1}(x_{1},x_{2})$ split into multiple context-dependent state representations, conditioned on $x_{2}$ value. Middle: ID test accuracy after standard training (100 epochs post training accuracy$>$0.99), showing Non-tree task significantly underperforms 2-Hop task ($|{\mathcal{X}}|=50$). Right: IICG heatmap from a model that achieved near-perfect ID accuracy (0.96) after extended training (36k epochs, $|{\mathcal{X}}|=50$, $N=50\text{k}$).
Many real-world reasoning tasks involve computational structures where a single variable affects the output through multiple paths, creating what we call path ambiguity. Our framework predicts that in such cases, Transformers struggle with forming unified representations of intermediate states that are theoretically equivalent, instead developing context-dependent state representations that vary based on the input context. In this section, we analyze Non-tree task (Fig. 2(d)) as a case study, where $x_{2}$ affects the output through two paths: as input to $f_{1}$ and directly to $f_{2}$. Following the definition of functional equivalence (Def. 3.1), this path ambiguity prevents the model from establishing the functional equivalence of two subsequences $(x_{1},x_{2})$ and $(x_{1}^{\prime},x_{2}^{\prime})$ that produce the same intermediate state $b$, unless they also share the same $x_{2}$ value ($x_{2}=x_{2}^{\prime}$). Consequently, the model creates context-dependent state representations rather than unifying them to represent true intermediate state $b$, as illustrated in Fig. 6 (Left).
In consequence, the path ambiguity hinders both generalization performance and efficiency, as the model must establish functional equivalence for each $x_{2}$-conditioned equivalent pair. Fig. 6 (Middle) shows that GPT-2 can fully generalize on ID test set of 2-Hop task within a reasonable time with increasing data size, but fails with Non-tree task, even provided with a near-exhaustive amount of possible ID combinations as training data. [^10] For $|\mathcal{X}|=50$ and $p_{\text{seen}}=0.7$, our largest run ($N=50\text{k}$) includes virtually the entire domain ($\approx 0.7^{2}\!\times|\mathcal{X}|^{3}\;\approx\;61\text{k}$ distinct ID triples). Notably, scaling to 1.5B parameters does not show significant improvement in the performance (Fig. 17 in App. G). Extremely prolonged training (36k epochs) with near-exhaustive ID combinations eventually achieves ID accuracy of 0.96. However, IICG analysis reveals no evidence of a unified intermediate state representation formation, with near-zero IICG scores when grouping by the intermediate state value $b$ (Fig. 6 (Right)). In contrast, grouping by $x_{2}$-conditioned intermediate state ($(b,x_{2})$) leads to high IICG scores, suggesting the formation of context-dependent state representations.
This context-dependence raises an interpretability concern, as standard linear probing-based techniques like logit lens [70, 71] would not reliably identify intermediate states. Moreover, our analysis provides insights into why LLMs fall short in planning tasks [29, 31, 30], as planning often requires correct tracking of intermediate states, which can influence outcomes through multiple paths.
5.5 CoT supervision improves data efficiency but still struggles with path ambiguity
Figure 7: Left: Power-law scaling of required dataset size vs. token set size for tasks with CoT supervision. $R^{2}>0.98$ for all linear fits. Middle: Comparison of ID test Accuracy of Non-tree task ($|{\mathcal{X}}|=50$) with and without CoT supervision (100 epochs post training accuracy$>$0.99). Right: IICG score comparison for Non-tree and 2-Hop task with CoT supervision ($|\mathcal{X}|=50$, $N=10\text{k}$). The scores are measured at each layer of intermediate state position $b$, based on two grouping strategies: $b$ and $(b,x_{2})$. Models are trained for 100 epochs after reaching training accuracy$>$0.99.
CoT supervision [27, 28] dramatically improves performance on multi-step reasoning tasks. We investigate how CoT interacts with the coverage principle and whether it can address the challenges observed in Sections 5.3 and 5.4. Specifically, we train models to sequentially generate intermediate states before final outputs, making 2-Hop a two-token prediction task: $(x_{1},x_{2},x_{3})\mapsto(b,t)$, for example. This explicit supervision substantially improves data efficiency (Fig. 7 (Left)), with the power-law exponent dropping from 2.58 to 1.76 in 3-Hop task, aligning with previous studies on the sample efficiency of CoT [79, 80, 81]. Remarkably, the scaling exponents measured for 2-Hop, 3-Hop, and even 5-Hop tasks become nearly identical with CoT supervision. We interpret this as CoT effectively ‘flattening’ multi-hop structures into sequences of single-hop tasks, reducing the compounding data requirements of deeper compositional structures.
However, we find models trained with CoT supervision still struggle with the path ambiguity in Non-tree task. Despite showing improvements with CoT supervision, the models fail to achieve perfect ID generalization under the same training conditions that yield perfect performance in 2-Hop task (Fig. 7 (Middle). IICG analysis (Right) reveals that the model’s representations remain partially context-dependent. For 2-Hop task, the representations cluster purely by intermediate states $b$, as indicated by the result that IICG measurement with $x_{2}$-conditioned states does not significantly shift the curve. In contrast, the IICG score for Non-tree task is significantly elevated at every layer with the same conditioning, suggesting the absence of disentangled state representation inside the model. We hypothesize this arises since CoT supervision does not give enough evidence that different $(x_{1},x_{2})$ pairs sharing the same $b$ should yield identical second-step outputs, as functional equivalence holds only when $x_{2}=x_{2}^{\prime}$. Hence, while CoT supervision helps with sequential computation by breaking down multi-hop structures, it may partially inherit the limitations on handling tasks with path ambiguities we describe in Sec. 5.4. Our analysis likely explains why LLMs struggle with complex planning tasks even when using CoT techniques and massive training data [82].
6 A taxonomy for understanding generalization mechanisms
Our coverage analysis reveals fundamental limits of pattern matching for compositional generalization. This raises a broader question: What other mechanisms allow neural networks to go beyond coverage boundaries? We sketch a preliminary taxonomy that distinguishes three complementary mechanisms and offers hypotheses about why models sometimes succeed beyond coverage [^11] The taxonomy is not exhaustive; Appendix H details limitations and open questions..
Type-I: Structure-based generalization occurs when models identify and utilize functionally equivalent components based on how primitive functions are composed. This is precisely what we formalize through the coverage principle: models learn that different input fragments yield identical results in shared contexts, enabling generalization to new fragment combinations. Crucially, this generalization remains bounded by coverage, and reliable generalization fails without sufficient functional equivalence evidence. Type-I describes the ceiling of pattern matching without explicit variable binding.
Type-II: Function property-based generalization exploits intrinsic properties of individual primitive functions (e.g., algebraic invariances such as commutativity), group-theoretic structure in modular arithmetic [63], or input irrelevance, where certain arguments never affect the output (e.g., $f(x_{1},x_{2})=f(x_{1})$ even when distractor $x_{2}$ is present [81]). Unlike Type-I, this can transcend coverage limitations by leveraging properties that hold across all inputs of a primitive. The Reversal Curse exemplifies the layered nature of compositional challenges across multiple generalization types. The coverage principle (Type-I) explains the fundamental failure: training on “$A$ is $B$” provides no functional equivalence evidence for “$B$ is $A^{-1}$”. Property-based approaches like bidirectional training partially succeed [83] by exploiting relationship invertibility (Type-II), using architectural modifications to learn inverse mappings from the same training data. However, recent evidence [42] shows that models still struggle once entities switch syntactic roles, indicating a remaining challenge in variable binding. [^12] See [42] for controlled concept–surface dissociation experiments.
Type-III: Shared-operator generalization emerges through reuse of identical primitive functions across computational positions (e.g., when $f_{1}=f_{2}$). Recurrent architectures [84] exemplify this through weight sharing across time steps, enabling processing of variable-length sequences [85]. In Transformers, inductive biases towards reuse of the same computation through parameter sharing [86, 87, 19] can improve generalization on compositional tasks beyond coverage. We interpret this mechanism as exploiting structural repetition without requiring Type-II’s algebraic property understanding.
Distinguishing mechanisms from phenomena
Prior categorizations focus on observed phenomena, for example, “systematic” versus “mix-and-match” [12]. We instead categorize the underlying mechanisms, as our coverage principle offers quantitative predictions for Type-I boundaries while identifying when Type-II or Type-III mechanisms become necessary. Many influential studies have examined tasks mixing functional equivalence, primitives’ intrinsic properties, and operator reuse within the same benchmark, making it difficult to pinpoint the true source of success or failure. We therefore advocate clearer experimental control and community discussion around this mechanistic distinction to sharpen future analyses of neural generalization.
Implications and future directions
Real compositional tasks typically involve combinations of all three types, making it difficult to isolate why models succeed or fail. We suggest diagnostic approaches based on the proposed taxonomy: coverage ablation tests for Type-I, algebraic identity probes for Type-II, and parameter-sharing ablations for Type-III. While preliminary, this taxonomy provides a conceptual scaffold for understanding the fundamental capabilities and limitations of neural networks on compositional reasoning. This work isolates and formalizes structure-based generalization (Type-I) to clarify its specific boundaries. When models succeed beyond our coverage predictions, e.g., through ignoring distractors without seeing all combinations or leveraging parameter sharing, we view these as exploiting Type-II or Type-III mechanisms rather than contradicting our framework. Our focus on Type-I suggests that fundamental systematicity challenges remain as long as models rely primarily on pattern matching, requiring architectural innovations that harness all three mechanisms and, crucially, incorporate explicit variable binding. We hope this preliminary taxonomy serves as a conversation starter, and confirming or refuting its utility is an empirical matter that we invite the community to explore.
7 Discussion
Systematicity without variable binding
The coverage principle pinpoints a fundamental question in the systematicity debate: How far can a neural network generalize compositionally when it relies only on pattern matching and has no mechanism for explicit variable binding? By formalizing the exact conditions under which pattern matching succeeds, we draw a principled boundary around what can be learned from input–output pairs alone. This boundary clarifies the limitations that Fodor and Pylyshyn [9] and Marcus [37] highlighted decades ago. Our results therefore complement approaches that encode compositional structure directly, whether through symbolic interfaces, slot-based representations, or neural modules [56, 57, 58, 59, 60, 88], and delineate what must still be overcome to reach truly systematic generalization.
Practical implications
Our framework provides valuable insights into various observed LLM behaviors. First, it accounts for the data-hungry nature of compositional tasks by demonstrating that robust coverage is required for reliable generalization [43]. Next, our framework explains why models struggle with generalizing long-tail knowledge [76, 77, 78], as low-frequency combinations naturally receive limited functional equivalence evidence. Similarly, failures in complex planning tasks [30, 82, 31, 89], even with CoT techniques, might be attributed to path ambiguities. Moreover, our coverage principle predicts the reversal curse phenomenon [16, 42, 83], as pattern-matching models fundamentally cannot generalize to reversed relationships without explicit functional equivalence evidence in training data.
Beyond explaining failures, our framework helps understand when and why standard interpretability techniques like linear probing or logit lens can fail. In addition, our analysis suggests targeted data augmentation strategies that maximize coverage by ensuring diverse shared contexts for functionally equivalent components. This can explain the success of strategic data augmentation methods [90, 91, 49, 48] and offers principled guidance. Finally, although our empirical study focuses on GPT-2–style Transformers, the coverage principle itself makes no architectural assumptions, and it applies to any learning system whose generalization mechanism is primarily pattern matching. Extending the analysis to recurrent (e.g., LSTM [84]), state-space [92], or convolutional [93] architectures, and more realistic data where multiple computational structures coexist within the same dataset [12, 15, 94] remains promising future work.
8 Conclusion
We introduce the coverage principle, a data-centric framework that specifies when pattern-matching learners can and cannot generalize compositionally. Our theoretical analysis and controlled experiments show that a Transformer’s success is tightly predicted by coverage: two-hop tasks obey a superquadratic data scaling law that does not improve with 20x parameter scaling, path ambiguity fragments internal representations, and CoT supervision helps only inside the coverage boundary.
These results isolate Type I structure-based generalization and reveal its limits. Our taxonomy then situates two additional mechanisms, Type II property exploitation and Type III shared-operator reuse, that help explain coverage “violations” reported in the literature. Yet all three mechanisms may still rely on sophisticated pattern matching rather than explicit variable binding, leaving the fundamental systematicity challenge unresolved.
Hence, the systematicity challenge posed by Fodor and Pylyshyn [9] and by Marcus [37] remains open. Matching human-like compositionality will likely require architectures that can bind and manipulate symbols independently of surface form, not merely scale up or refine current pattern-matching strategies. We hope the coverage principle and the accompanying taxonomy provide a clear target and a roadmap for the next stage of innovation.
Acknowledgments and Disclosure of Funding
We would like to thank Jaewon Oh, Jihoon Ha, Hyowon Cho, Yunjae Won, Seongyun Lee, and Boshi Wang for their valuable feedback on our work.
Appendix A Limitations
We deliberately restrict to synthetic tasks to isolate structure-based limits without confounds from lexical or domain priors. We leave extending the coverage analysis to natural data as a future work. Additionally, our experiments focus on autoregressive Transformer architectures, and the applicability of the coverage principle to other architectures (e.g., state-space models) remains to be validated.
Appendix B Detailed experimental setup
B.1 Dataset construction details
We now provide detailed information about our dataset construction process. While we primarily explain this process for the 2-Hop task, we follow similar procedures for the other compositional structures.
Vocabulary and Token Representation
For a task with token set size $|\mathcal{X}|$, we create $|\mathcal{X}|$ special tokens of the form <t_0>, <t_1>, $\ldots$, <t_$(|\mathcal{X}|-1)$>, which we append to the standard GPT-2 vocabulary. We also add special tokens </a> to mark the end of sequences. For Chain-of-Thought (CoT) experiments, intermediate computations are represented in the target sequence as the actual intermediate token.
Function Construction
For the 2-Hop task, we construct two primitive functions $f_{1}:\mathcal{X}^{2}\rightarrow\mathcal{X}$ and $f_{2}:\mathcal{X}^{2}\rightarrow\mathcal{X}$ by randomly mapping from their respective domains to the codomain $\mathcal{X}$. We create the domain by taking the Cartesian product of the token set with itself. For each function, we randomly designate a fraction $p_{\text{seen}}=0.7$ of its domain as the "seen" portion, resulting in sets $S_{f_{1}}$ and $S_{f_{2}}$.
Dataset Generation Algorithm
To generate the training dataset, we first identify all possible combinations where both primitive operations come from their respective "seen" domains. Specifically, we find all valid tuples $(x_{1},x_{2},x_{3},t)$ such that:
$(x_{1},x_{2})$ $\in\text{domain}(S_{f_{1}})$ (1)
$(f_{1}(x_{1},x_{2}),x_{3})$ $\in\text{domain}(S_{f_{2}})$ (2)
$t$ $=f_{2}(f_{1}(x_{1},x_{2}),x_{3})$ (3)
From this set of all possible in-domain combinations, we uniformly sample $N$ examples to form our training dataset. When the number of possible combinations exceeds $N$, this sampling ensures the model sees only a subset of possible in-domain combinations.
Test Set Construction
We carefully construct test sets to evaluate the model’s generalization capabilities across different coverage conditions. Our test sets contain:
  • In-Domain (ID) Test Set: Combinations not seen during training but where both primitive operations were observed in other contexts. These examples may lie within the coverage as defined by our framework.
  • Out-of-Domain (OOD) Test Set: Examples where at least one primitive operation was never observed in training. These fall outside the coverage boundary.
Input-Output Format
The dataset is formatted for auto-regressive token prediction. For the standard 2-Hop task, inputs comprise three tokens representing $x_{1}$, $x_{2}$, and $x_{3}$, while the target includes these input tokens followed by the prediction $t$ and an end marker. Below are the examples of the dataset format for different settings.
  • Standard Format:
    • Input: <t_5><t_12><t_3>
    • Target Completion: <t_17></a>
    • The model must predict the final output token followed by the end marker.
  • Chain-of-Thought Format:
    • Input: <t_5><t_12><t_3>
    • Target Completion: <t_9><t_17></a>
    • The model must first predict the intermediate computation result <t_9> (where <t_9> = $f_{1}$(<t_5>, <t_12>)), followed by the final output.
  • Partial Computation Format ($f_{1}$):
    • Input: <t_5><t_12>
    • Target Completion: <t_9></a>
    • These examples represent the primitive function applications used to construct the full compositional task.
For the other compositional tasks, we follow analogous construction procedures, adjusting the number of input tokens and the composition structure based on the specific task’s requirements. For example, Parallel 2-Hop requires four input tokens, while 3-Hop follows a three-step composition chain requiring appropriate modifications to the function construction and sampling procedures.
B.2 Training details
Table 1: Model configurations for different GPT-2 variants used in our experiments
Configuration GPT-2-Small GPT-2 GPT-2-XL
Number of Attention Heads 6 12 25
Number of Layers 4 8 48
Hidden Dimension 768 768 1600
Total Parameters 68M 96M 1.5B
For our experiments, we employ three GPT-2 model variants of increasing size: GPT-2-Small (68M parameters), GPT-2 (96M parameters), and GPT-2-XL (1.5B parameters). As shown in Tab. 1, GPT-2-Small consists of 4 layers with 6 attention heads and a hidden dimension of 768. The standard GPT-2 configuration used in most experiments features 8 layers with 12 attention heads while maintaining the same hidden dimension of 768. Our largest model, GPT-2-XL, significantly scales up the architecture with 48 layers, 25 attention heads, and an increased hidden dimension of 1600. The implementation follows the codebase from [19].
We train all models using the AdamW optimizer with beta values of (0.9, 0.999) and epsilon of 1e-8. We set the learning rate to 8e-4 with a weight decay of 0.1. A batch size of 16,384 is used, with full gradient descent applied for datasets smaller than the batch size. All training is conducted with mixed precision (fp16) on 4 NVIDIA A100 GPUs with 80GB memory each. We employ a constant learning rate schedule with a linear warmup period of 2,000 steps. This standardized training configuration is maintained across all experiments to ensure fair comparisons between different task structures and dataset sizes, unless explicitly varied in specific ablation studies.
Appendix C Implementation details for the coverage determination algorithm
Input: Training examples $\mathcal{D}=\{(x_{i},f(x_{i}))\}$, where $x_{i}\in\mathcal{X}^{n}$ and $f(x_{i})\in\mathcal{X}$
Minimum evidence threshold $k\geq 1$
Output: Coverage set $\text{Cover}(\mathcal{D})$
/* STEP 1: Build behavior maps for each subsequence pattern */
foreach subset $I\subset[n]$, $I\neq\emptyset$, $I\neq[n]$ do
       $\text{Behavior}_{I}\leftarrow$ map from subsequence $x_{I}$ to the mapping $\{x_{[n]\setminus I}\mapsto f(x)\mid x\in\mathcal{D}\}$
end foreach
/* STEP 2: Identify functionally equivalent subsequences */
foreach subset $I\subset[n]$, $I\neq\emptyset$, $I\neq[n]$ do
       $UF_{I}\leftarrow\text{new UnionFind}()$
      foreach pair of subsequences $(\alpha,\beta)$ in $\text{Behavior}_{I}$ do
[…]
            if No contradictions in SharedComplements and matching evidence $\geq k$ then
                   $UF_{I}.\text{Union}(\alpha,\beta)$ ;
                    // Mark as functionally equivalent
             end if
       end foreach
      $\text{EquivClasses}_{I}\leftarrow UF_{I}$
end foreach
/* STEP 3: Build substitution graph */
$G\leftarrow\text{empty graph}$ with nodes for all $x\in\mathcal{X}^{n}$
foreach pair of inputs $(x,y)$ with $f(x)=f(y)$ do
       foreach subset $I$ where $x$ and $y$ differ only on indices in $I$ do
             if $\text{EquivClasses}_{I}.\text{Find}(x_{I})=\text{EquivClasses}_{I}.\text{Find}(y_{I})$ then
                   Add edge $(x,y)$ to $G$
                  break
             end if
       end foreach
end foreach
/* STEP 4: Determine coverage */
$\text{Cover}(\mathcal{D})\leftarrow\bigcup_{x\in\mathcal{D}}$ ConnectedComponent$(G,x)$
return $\text{Cover}(\mathcal{D})$
Algorithm 1 $k$-Coverage Determination Algorithm
Algorithm 1 presents our approach to computing the coverage set with a minimum evidence threshold $k$. The algorithm works in four main stages:
Stage 1: Behavior mapping
We first analyze the training data to create a mapping of behaviors for each possible subsequence of the input. For each subset of indices $I$, we record how different subsequences $x_{I}$ behave when paired with their complements $x_{[n]\setminus I}$, essentially mapping each subsequence to a function from complements to outputs.
Stage 2: Equivalence class construction
For each subset of indices $I$, we build equivalence classes of subsequences that exhibit functionally identical behavior. Two subsequences are considered equivalent only if: (1) they share at least $k$ distinct complements where they produce the same output, and (2) they never produce different outputs when given the same complement (no contradictions). We use a Union-Find data structure to efficiently track and merge these equivalence classes. The Union-Find (or Disjoint-Set) data structure efficiently maintains a collection of disjoint sets, supporting two key operations: (1) Find - determine which set an element belongs to, and (2) Union - merge two sets.
Stage 3: Substitution Graph Construction
We construct a graph where nodes represent input sequences from our training and test sets, rather than the entire domain space (which would be computationally prohibitive for large token sets). We add an edge between two inputs $x$ and $y$ if and only if: (1) they produce the same output, (2) they differ only in one subsequence position set $I$, and (3) their differing subsequences belong to the same equivalence class. This graph represents the space of safe substitutions where one can replace a subsequence with a functionally equivalent alternative without changing the expected output. Our implementation uses parallel processing to efficiently construct this graph even for large datasets.
Stage 4: Coverage computation
Finally, we compute the coverage set by taking the union of all connected components in the substitution graph that contain at least one training example. This set comprises all inputs that are reachable from the training data through chains of equivalent subsequence substitutions.
Appendix D Detailed analysis for representation unification experiments
D.1 Causal Tracing Methodology
To analyze the causal role of specific hidden representations in our Transformer model, we employ causal tracing, a technique that measures the effect of intervening on intermediate activations during inference [67, 66]. Specifically, we measure the causal effect using the indirect effect metric defined in [95]. This methodology allows us to identify which components and positions in the model most strongly contribute to compositional generalization. We illustrate the measurement with 2-Hop task.
Our analysis begins by collecting three types of computational traces:
  1. 1.
    Clean run ($G$): We run the model on a compositional task with input $(x_{1},x_{2},x_{3})$ where the corresponding output is $t=f_{2}(f_{1}(x_{1},x_{2}),x_{3})$.
  2. 2.
    Corrupted run ($G^{*}$): We replace the original input with a corrupted version by changing the first two tokens $(x_{1},x_{2})$ to $(x_{1}^{\prime},x_{2}^{\prime})$, where $f_{1}(x_{1}^{\prime},x_{2}^{\prime})\neq f_{1}(x_{1},x_{2})$. This ensures that the model produces a different final output $t^{*}\neq t$. During this run, we cache all hidden states ${h^{*}}_{i}^{(\ell)}$ for each token position $i$ and layer $\ell$.
  3. 3.
    Patched run ($G[\leftarrow{h^{*}}_{i}^{(\ell)}]$): We run the model on the input from the clean run, but at a specific token position $i$ and layer $\ell$, we replace the hidden state with the corresponding state from the corrupted run.
To quantify the causal effect of a specific hidden state $h_{i}^{(\ell)}$ on the model’s prediction, we measure the Indirect Effect (IE):
$\text{IE}_{h_{i}^{(\ell)}}=\frac{p[\leftarrow{h^{*}}_{i}^{(\ell)}](t^{*})-p(t^{*})}{p^{*}(t^{*})-p(t^{*})}$ (4)
where:
  • $p(t^{*})$ is the probability assigned to the corrupted output $t^{*}$ in the clean run $G$
  • $p^{*}(t^{*})$ is the probability assigned to the corrupted output $t^{*}$ in the corrupted run $G^{*}$
  • $p[\leftarrow{h^{*}}_{i}^{(\ell)}](t^{*})$ is the probability assigned to the corrupted output $t^{*}$ in the patched run $G[\leftarrow{h^{*}}_{i}^{(\ell)}]$
This metric quantifies how much corruption in a particular state affects the overall outcome. An IE value close to 1 indicates that the corruption of the state $h_{i}^{(\ell)}$ to ${h^{*}}_{i}^{(\ell)}$ alone almost completely changes the prediction to that of the corrupted run, suggesting that this state is causally important for the computation. Conversely, an IE value close to 0 indicates that the state has minimal causal impact on the prediction.
In our experiments, we apply causal tracing to analyze different subsets of test data categorized by their $k$-cutoff values, where $k$ represents the minimum evidence threshold required for functional equivalence (as defined in Sec. 3 of the main text). This allows us to correlate the strength of functional equivalence evidence with the formation of unified internal representations.
D.2 Causal tracing results for each $k$-cutoff value in 2-Hop task
Figure 8 displays the causal tracing results for the 2-Hop task, broken down by different $k$-cutoff values. We observe that the causal patterns are similar across different $k$-cutoff values, with slight differences in where and how strongly the causal effects manifest in the model. This suggests that once an example falls within coverage (even with minimal evidence, $k=1$), the model forms internal representations that play similar causal roles in prediction.
Figure 8: Causal tracing results for the 2-Hop task across different $k$-cutoff values, showing Indirect Effect (IE) scores at each layer and position.
D.3 Token set size ablation
We show that the observed patterns of cosine similarity analysis and causal tracing in the 2-Hop task are consistent across different token set sizes $|{\mathcal{X}}|$. For $|{\mathcal{X}}|=70,100,150,200$, we analyze model checkpoints with training dataset size $N=\hat{N}_{\text{req}}(|{\mathcal{X}}|)$ that achieve training accuracy > 0.99. Figure Fig. 9 shows the results, indicating strong representation clustering at the lower layers of position $x_{2}$ for all cases. The causal tracing results in Fig. 10 show that the clustered functional equivalence representations at the lower layers of position $x_{2}$ play a causal role in determining the model’s final prediction.
Figure 9: IICG heatmap across different token set sizes, showing consistent representation clustering patterns.
Figure 10: Causal tracing results showing indirect effect heatmaps for different token set sizes $|{\mathcal{X}}|$.
D.4 Task ablation
We show that GPT-2 models trained on Parallel-2-Hop and 3-Hop tasks exhibit the same patterns: clustered functional equivalence representations of intermediate states at specific layers and positions, confirmed through cosine similarity analysis, with causal tracing analysis verifying their role in model predictions. For both tasks, we analyze with $|{\mathcal{X}}|=50$ and examine model checkpoints with training dataset size $N=\hat{N}_{\text{req}}(|{\mathcal{X}}|)$ that achieve training accuracy > 0.99.
Figures 11 and 12 show the results for the Parallel-2-Hop task. The IICG patterns reveal strong representation clustering at mid-layers: at positions $x_{2}$ and $x_{3}$ when grouped by $b_{1}=f_{1}(x_{1},x_{2})$, and at position $x_{4}$ when grouped by $b_{2}=f_{2}(x_{3},x_{4})$. Causal tracing confirms the causal role of these clustered representations in the model’s predictions.
Figure 11: IICG heatmap for Parallel-2-Hop task with grouping strategies based on $b_{1}=f_{1}(x_{1},x_{2})$ (Left), $b_{2}=f_{2}(x_{3},x_{4})$ (Middle), and $t=f_{3}(b_{1},b_{2})$ (Right).
Figure 12: Causal tracing results showing indirect effect heatmap for Parallel-2-Hop task. Left: perturbation with different $(x_{1},x_{2})$ pair leading to a different $b_{1}$ value. Right: perturbation with different $(x_{3},x_{4})$ pair leading to a different $b_{2}$ value.
Similarly, Figures 13 and 14 show results for the 3-Hop task. The IICG patterns exhibit strong representation clustering at mid-layers: at position $x_{3}$ when grouped by $b_{1}=f_{1}(x_{1},x_{2})$, and at position $x_{3}$ when grouped by $b_{2}=f_{2}(b_{1},x_{3})$. Causal tracing again confirms the causal importance of these representations.
Figure 13: IICG heatmap for 3-Hop task with grouping strategies based on $b_{1}=f_{1}(x_{1},x_{2})$ (Left), $b_{2}=f_{2}(b_{1},x_{3})$ (Middle), and $t=f_{3}(b_{2},x_{4})$ (Right).
Figure 14: Causal tracing results showing indirect effect heatmap for 3-Hop task. Left: perturbation with different $(x_{1},x_{2})$ pair leading to different $b_{1}$ value. Right: perturbation leading to different $b_{2}=f_{2}(b_{1},x_{3})$ value.
These results demonstrate that the formation of clustered intermediate state representations and their causal role in compositional generalization is a consistent phenomenon across different task structures, supporting the generality of our findings beyond the 2-Hop task.
Appendix E Derivation of Result 5.1 (power–law lower bound)
Problem setting
Let ${\mathcal{X}}$ be a finite token set with cardinality $|{\mathcal{X}}|$. The target mapping
$f:{\mathcal{X}}^{3}\longrightarrow{\mathcal{X}},\qquad f(x_{1},x_{2},x_{3})=f_{2}\bigl{(}f_{1}(x_{1},x_{2}),\,x_{3}\bigr{)}$
is a two-hop composition of unknown primitives $f_{1}:{\mathcal{X}}^{2}\to{\mathcal{X}}$ and $f_{2}:{\mathcal{X}}^{2}\to{\mathcal{X}}$. Write $b=f_{1}(x_{1},x_{2})\in{\mathcal{X}}$ for the intermediate state. Throughout the proof we impose two assumptions.
  1. (A1)
    Balanced classes: Each intermediate value $b\in{\mathcal{X}}$ is realized by exactly $|{\mathcal{X}}|$ first-hop pairs, i.e., the sets $E_{b}:=\{(x_{1},x_{2})\in{\mathcal{X}}^{2}:f_{1}(x_{1},x_{2})=b\}$ all have size $|{\mathcal{X}}|$ and form a partition of ${\mathcal{X}}^{2}$.
  2. (A2)
    Uniform training sampler: The training set $D$ contains $N$ triples drawn uniformly with replacement from the domain ${\mathcal{X}}^{3}$.
Functional $k$-equivalence and learner model
For two first-hop fragments $a,a^{\prime}\in{\mathcal{X}}^{2}$ define
$a\sim a^{\prime}\quad:\Longleftrightarrow\quad f_{1}(a)=f_{1}(a^{\prime}).$
They are functionally $k$-equivalent w.r.t. $D$ (denoted $a\equiv^{k}_{D}a^{\prime}$) when there exist $k$ distinct contexts $c_{1},\dots,c_{k}\in{\mathcal{X}}$ such that for every $r\leq k$ both $(a,c_{r})$ and $(a^{\prime},c_{r})$ appear in $D$ and $f(a,c_{r})=f(a^{\prime},c_{r})$. We call each $(a,c_{r}),(a^{\prime},c_{r})$ an evidence pair.
The coverage principle by itself is only a necessity statement: outside the $k$-coverage region a purely pattern matching learner’s predictions are unconstrained. To convert this into a sufficient data condition we adopt an explicit inductive bias, matching the premise of Result 5.1:
Learner assumption: Whenever two fragments become linked by $k$ independent (i.e., pairwise from distinct contexts) evidence pairs, the learner treats them as functionally equivalent; the learner also propagates equivalence transitively along chains of such links.
With this rule in place the relevant structure inside each class $E_{b}$ is the $k$-evidence graph: vertices are the $|{\mathcal{X}}|$ first-hop pairs in the class and an edge connects two vertices whenever the pair is observed at least $k$ times in shared contexts. If that graph is connected (every vertex reachable from every other), then every fragment in $E_{b}$ is linked by a chain of $k$-evidence steps and is therefore recognized as equivalent by the learner. Hence
Data sufficiency criterion: If the $k$-evidence graph of each class $E_{b}$ is connected, the learner generalizes perfectly to all in-domain (ID) inputs.
This criterion requires connectivity, and we aim to derive the condition to yield the connectivity with high probability using Erdős–Rényi model [96, 97]. The minimal dataset size achieving this with high probability is denoted $N_{\mathrm{req}}(|{\mathcal{X}}|,k)$.
Note that the $k$-evidence graph on $E_{b}$ is the restriction of the substitution graph $\mathcal{G}_{D,k}$ (Def. 3.2) to the vertex set $E_{b}\times\{x_{3}\}$ for any fixed $x_{3}$. Connectivity of every class therefore implies that every first-hop fragment lies in the same connected component as some training input, i.e., the entire in-domain set is contained in $k$-coverage. Under the learner assumption, this is both necessary and sufficient for perfect ID generalization, yielding Res. 5.1.
E.1 Step 1: probability of a single evidence pair
Evidence pair probability
Fix two distinct first–hop fragments $i,j\in E_{b}$ and a context $c\in{\mathcal{X}}$. We want to find $p_{1}$, the probability that context $c$ provides an evidence pair for the functional equivalence of fragments $i$ and $j$:
$p_{1}:=\Pr[(i,c)\in D\text{ and }(j,c)\in D]$
Let $q:=1/|{\mathcal{X}}|^{3}$ denote the probability of drawing any specific triple in a single draw. Using the inclusion–exclusion principle:
$p_{1}$ $=\Pr[(i,c)\in D\text{ and }(j,c)\in D]$
$=1-\Pr[(i,c)\not\in D]-\Pr[(j,c)\not\in D]+\Pr[(i,c)\not\in D\text{ and }(j,c)\not\in D]$
$=1-(1-q)^{N}-(1-q)^{N}+(1-2q)^{N}$ (S1.1)
For $q\ll 1$ (which holds when $|{\mathcal{X}}|$ is large), we can use Taylor expansion:
$(1-q)^{N}$ $=1-Nq+\frac{N(N-1)}{2}q^{2}+O(q^{3})$
$(1-2q)^{N}$ $=1-2Nq+\frac{N(N-1)}{2}(2q)^{2}+O(q^{3})$ (S1.2)
Substituting these approximations:
$p_{1}$ $=1-2\left(1-Nq+\frac{N(N-1)}{2}q^{2}\right)+\left(1-2Nq+2N(N-1)q^{2}\right)+O(q^{3})$
$=1-2+2Nq-N(N-1)q^{2}+1-2Nq+2N(N-1)q^{2}+O(q^{3})$
$=N(N-1)q^{2}+O(q^{3})$
$=\frac{N(N-1)}{|{\mathcal{X}}|^{6}}+O\left(\frac{1}{|{\mathcal{X}}|^{9}}\right)$
$=\frac{N^{2}}{|{\mathcal{X}}|^{6}}(1+o(1))$ (S1.3)
Therefore, in the regime of interest ($N\gg|{\mathcal{X}}|$ but $N\ll|{\mathcal{X}}|^{3}$):
$\boxed{p_{1}=\frac{N^{2}}{|{\mathcal{X}}|^{6}}(1+o(1))}$
Equation (E.1) gives an exact expression for the probability (up to lower-order terms) that the single context $c$ provides an evidence pair for the functional equivalence of fragments $i$ and $j$.
Remark on “lucky coincidences”
Because functional $k$-equivalence demands consistency across all $k$ evidences, a single coincidental equality $f(i,c)=f(j,c)$ with $i\not\sim j$ can only masquerade as evidence when $k=1$ and the dataset lacks any contradicting context $c^{\prime}$. For $k\geq 2$ the joint probability that two independent contexts simultaneously produce such coincidences is $|{\mathcal{X}}|^{-k}$ per fragment pair and hence negligible relative to the isolate probability once $N=\Omega(|{\mathcal{X}}|^{2})$. Consequently, the effects of such “lucky coincidences” on the lower bound in Res. 5.1 can be neglected.
E.2 Step 2: probability of observing $k$ evidences for one fixed pair
Having established the probability $p_{1}$ for a single evidence pair in Step 1, we now derive $p_{k}$, the probability that a dataset provides at least $k$ distinct contexts as evidence for the functional equivalence of two fragments $i$ and $j$. This probability will determine the edge probability in the random graph model analyzed in Step 3.
Indicators for one fragment pair
Fix two distinct first-hop fragments $i,j\in E_{b}$ and, for each context (third token) $c\in{\mathcal{X}}$, set
[…]
Thus $Z_{ij}(c)=1$ exactly when the single context $c$ supplies an evidence pair for the functional equivalence of $i$ and $j$.
Single-context success probability
From (E.1),
$p_{1}:=\Pr[Z_{ij}(c)=1]=\frac{N^{2}}{|{\mathcal{X}}|^{6}}\bigl{(}1+o(1)\bigr{)},\qquad(|{\mathcal{X}}|\to\infty).$
Negatively correlated counts and an i.i.d. surrogate
Because all $N$ draws come from a single multinomial $\bigl{(}N,1/|{\mathcal{X}}|^{3}\bigr{)}$, the indicators $\{Z_{ij}(c)\}_{c\in{\mathcal{X}}}$ are negatively correlated: drawing many triples with one context leaves fewer draws for the others. Negative correlation decreases the probability that several $Z_{ij}(c)$ equal 1 simultaneously. Hence the tail probability for the true count
$Y_{ij}:=\sum_{c\in{\mathcal{X}}}Z_{ij}(c)$
is upper-bounded by the tail of an i.i.d. binomial variable with the same single-trial success probability $p_{1}$. Concretely, define
$Y_{ij}^{\star}\sim\mathrm{Binom}\bigl{(}|{\mathcal{X}}|,p_{1}\bigr{)},\qquad p_{1}=\frac{N^{2}}{|{\mathcal{X}}|^{6}}\bigl{(}1+o(1)\bigr{)}.$
Then for every real $t$
$\Pr\bigl{[}Y_{ij}\geq t\bigr{]}\leq\Pr\bigl{[}Y_{ij}^{\star}\geq t\bigr{]}.$
Using $Y_{ij}^{\star}$ therefore overestimates the chance of obtaining $k$ or more distinct evidence contexts, which is conservative for our goal of deriving a lower bound on the required dataset size $N$.
Since $Y_{ij}^{\star}$ is binomial with mean
$\mu:=\mathbb{E}[Y_{ij}^{\star}]=|{\mathcal{X}}|\,p_{1}=\frac{N^{2}}{|{\mathcal{X}}|^{5}}\bigl{(}1+o(1)\bigr{)},$
we may work henceforth with $Y_{ij}^{\star}$ alone; the resulting bounds apply verbatim to the original $Y_{ij}$.
Poisson tail via Le Cam
[…]
$\Pr[Y_{ij}\geq k]\leq\Pr[Y_{ij}^{\star}\geq k]=\sum_{r=k}^{\infty}\frac{e^{-\mu}\mu^{r}}{r!}\leq\frac{\mu^{k}}{k!}\bigl{(}1+o(1)\bigr{)}.$
This upper bound $p_{k}:=\mu^{k}/k!(1+o(1))$ will be the edge probability used in the connectivity threshold of Step 3.
E.3 Step 3: connectivity inside each equivalence class
With the edge probability $p_{k}$ from Step 2, we now analyze when the $k$-evidence graphs become connected. Recall that under our learner assumption, perfect generalization requires every equivalence class to form a connected component in the $k$-evidence graph. We model this as a random graph connectivity problem.
Random graph construction
Fix one balanced class $E_{b}$ of size $n:=|{\mathcal{X}}|$ (Assumption (A1)). Create a graph $G_{b}$ whose vertices are the first-hop pairs in $E_{b}$, and place an undirected edge $\{i,j\}$ exactly when the pair $(i,j)$ has been observed in at least $k$ distinct shared contexts. By Step E.2, each potential edge appears with probability
$p_{k}=\Pr\bigl{[}Y_{ij}\geq k\bigr{]}=\frac{\mu^{k}}{k!}\,,\qquad\mu=\frac{N^{2}}{|{\mathcal{X}}|^{5}}.$
Why $G_{b}$ can be approximated as Erdős–Rényi
As mentioned earlier, the dependence among distinct edges in $G_{b}$ arises from the constraint that the total sample size is $N$, which induces a negative correlation. Such negative dependence reduces the likelihood of simultaneously creating many edges. Consequently, viewing $G_{b}$ as an independent Erdős–Rényi graph $G(n,p_{k})$ provides a conservative model, and any threshold we derive for connectivity under independence remains valid (or becomes easier to satisfy).
Classwise connectivity threshold
For Erdős–Rényi graphs, the classical result of Erdős–Rényi [97] states
$\Pr\bigl{[}G(n,p)\text{ is connected}\bigr{]}\xrightarrow[n\to\infty]{}1\quad\Longleftrightarrow\quad p\geq\frac{\log n+\omega(1)}{n}.$
Setting $n=|{\mathcal{X}}|$ and $p=p_{k}$ yields the requirement
$p_{k}=\frac{\mu^{k}}{k!}\geq\frac{\log|{\mathcal{X}}|}{|{\mathcal{X}}|}(1+o(1)).$
Substituting $\mu=N^{2}/|{\mathcal{X}}|^{5}$ and rearranging gives
$\frac{N^{2k}}{|{\mathcal{X}}|^{5k}\,k!}\gtrsim\frac{\log|{\mathcal{X}}|}{|{\mathcal{X}}|},\qquad\Longrightarrow\qquad N\gtrsim|{\mathcal{X}}|^{\frac{5k-1}{2k}}\,\bigl{(}k!\log|{\mathcal{X}}|\bigr{)}^{1/(2k)}.$
Because $(k!\log|{\mathcal{X}}|)^{1/(2k)}=(\log|{\mathcal{X}}|)^{O(1/k)}$ grows only poly-logarithmically, we hide it inside a $\tilde{\Omega}(\,\cdot\,)$:
[…]
Since every class must be connected to achieve the data–sufficiency criterion in the problem setting, we conclude
[…]
Appendix F Additional results for power-law scaling analysis
F.1 Measurement protocol for $N_{\mathrm{req}}$
To empirically determine the minimum dataset size required for reliable compositional generalization ($N_{\mathrm{req}}$), we develop a measurement protocol that accounts for practical computational constraints while ensuring robustness. For each token set size $|\mathcal{X}|$ and task structure, we test multiple dataset sizes until we identify the threshold point where the model successfully generalizes to the ID test set.
Specifically, our criterion for “reliable generalization” on ID is defined as:
  • The model must reach ID test accuracy of 0.99 within 100 epochs after achieving training accuracy > 0.99.
This protocol balances several considerations:
  1. 1.
    Training-to-generalization delay: Larger datasets naturally require more iterations to fit training data. By measuring epochs after reaching training accuracy > 0.99, we focus on the generalization gap rather than conflating it with initial training difficulty.
  2. 2.
    Epoch-based measurement: Using epochs rather than raw training steps ensures that the model sees each functional equivalence evidence approximately the same number of times, regardless of dataset size. This provides a fairer comparison across different dataset sizes.
  3. 3.
    Practical time constraints: While indefinite training might eventually yield generalization with smaller datasets, we established a reasonable upper bound (100 epochs post-training convergence) to reflect practical limitations.
  4. 4.
    Measurement precision: For each identified $N_{\mathrm{req}}$, we verified that 75% of this dataset size consistently failed to meet our generalization criterion. This establishes that our measurement error is at most $-\log(0.75)=0.125$ in log scale, providing confidence in the derived power-law exponents.
F.2 Measured power-law scaling constants across task structures and model sizes
Using our measurement protocol, we measure the required dataset size $N_{\mathrm{req}}$ across three different compositional structures (2-Hop, Parallel-2-Hop, and 3-Hop) and three model scales (68M, 96M, and 1.5B parameters). For each task structure, we vary the token set size $|\mathcal{X}|$ from 50 to 200, allowing us to observe the scaling relationship.
Table 2 presents the power-law exponents obtained by linear fitting $\log(|\mathcal{X}|)$ vs. $\log(N_{\mathrm{req}})$ plots, all with $R^{2}>0.99$. The consistency of exponents across model sizes suggests that the observed power-law scaling relates to properties of the compositional tasks themselves, rather than model capacity. This observation aligns with our theoretical derivation in Section 5.1, which predicts that the required dataset size scales at least quadratically with token set size.
Table 2: Power law exponents for different tasks and GPT-2 sizes, obtained by linear fitting $\log(|\mathcal{X}|)$ vs. $\log(N_{\mathrm{req}})$ plots. $R^{2}>0.99$ for all linear fitting.
Model Size 2-hop Parallel-2-hop 3-hop
68M 2.13 2.47 2.61
96M 2.26 2.35 2.50
1.5B 2.28 2.17 2.60
F.3 Robustness to hyperparameter variations
Figure 15: Robustness of power-law scaling relationship to hyperparameter variations in the 2-Hop task with $|\mathcal{X}|=50$. Each line shows the training and test accuracy curves for a different configuration: (1) baseline, (2) reduced learning rate (4e-4, half of baseline), (3) reduced weight decay (0.01, one-tenth of baseline), and (4) changed generalization criteria (test accuracy > 0.95 within 10 epochs after training accuracy > 0.95). $R^{2}>0.99$ for all linear fitting.
To verify that our observed power-law scaling relationship is not an artifact of specific hyperparameter choices, we conduct ablation studies with modified training configurations. Figure 15 demonstrates that for the 2-Hop task with $|\mathcal{X}|=50$, the following changes did not significantly affect the measured $N_{\mathrm{req}}$ or the derived power-law exponent:
  1. 1.
    Learning rate reduction: Halving the learning rate from 8e-4 to 4e-4
  2. 2.
    Weight decay reduction: Decreasing weight decay by a factor of 10 (from 0.1 to 0.01)
  3. 3.
    Generalization criteria modification: Requiring test accuracy > 0.95 within 10 epochs after training accuracy > 0.95
This robustness to hyperparameter variations suggests that the power-law relationship between token set size and required dataset size is primarily a property of the compositional generalization process, rather than an artifact of specific optimization settings.
Appendix G Detailed analysis for Non-tree task
This section provides additional analyses that support our findings in Sec. 5.4 regarding the challenges of path ambiguity in the Non-tree task.
G.1 Coverage analysis
Figure 16: Coverage analysis for Non-tree task with $|\mathcal{X}|=50$. The graph shows the percentage of ID test data covered at different $k$ values across various dataset sizes ($N$). Compared to the 2-Hop task (Fig. 3, left), Non-tree has significantly lower coverage at equivalent dataset sizes, indicating that path ambiguity impedes the formation of functional equivalence relationships.
Fig. 16 demonstrates that with equivalent training dataset sizes, a smaller percentage of ID test examples fall inside $k$-coverage for the Non-tree task compared to the 2-Hop task shown in Fig. 3 (Left). This aligns with our theoretical analysis in Sec. 5.4, which predicts that path ambiguity limits the establishment of functional equivalence relationships between input subsequences, as the model cannot generalize across different $x_{2}$ values in the Non-tree structure even when they produce the same intermediate state $b=f_{1}(x_{1},x_{2})$.
G.2 Effect of model scaling
Figure 17: ID test accuracy comparison between GPT-2 (96M parameters) and GPT-2-XL (1.5B parameters) on the Non-tree task with $|\mathcal{X}|=50$, measured 100 epochs after training accuracy exceeds 0.99. Despite the 15x increase in parameter count, the accuracy does not increase.
Fig. 17 shows that scaling up the model size to GPT-2-XL (1.5B parameters) does not significantly improve generalization performance on the Non-tree task, even when measured 100 epochs after reaching training accuracy > 0.99. This suggests that the challenges posed by path ambiguity cannot be overcome simply by increasing model capacity, supporting our claim that the limitation is structural rather than related to model capacity.
G.3 Representation analysis in successful generalization
For a model that eventually achieved near-perfect ID accuracy (0.96) after extended training (36k epochs, $|\mathcal{X}|=50$, $N=50\text{k}$), we conduct causal tracing analysis to understand how it achieves generalization despite path ambiguity.
Figure 18: Causal tracing analysis for the Non-tree model after extended training. The heatmap shows indirect effect values across different layer-token positions. Left: perturbation leading to different intermediate state $b=f_{1}(x_{1},x_{2})$. Middle: same $b$ but different $x_{2}$. Right: different $b$ and $x_{2}$.
The causal tracing results in Fig. 18 reveal how the model achieves generalization in the presence of path ambiguity. Across all perturbation strategies, the model’s predictions show strong causal dependence on representations at both the $x_{1}$ and $x_{2}$ positions, indicating reliance on direct access to both input tokens rather than an abstracted intermediate computation. This pattern contrasts sharply with the 2-Hop task, where causal effects concentrate primarily at positions corresponding to clustered functional equivalence representations.
This analysis demonstrates that even models achieving high accuracy on Non-tree tasks do so by developing context-dependent representations rather than unified abstractions of intermediate states. The model forms separate computational pathways conditioned on the $x_{2}$ value, rather than learning a single unified representation of the intermediate state $b=f_{1}(x_{1},x_{2})$. This represents a fundamentally different solution strategy compared to the 2-Hop task, with implications for both generalization capability and interpretability.
Appendix H Detailed discussion on the taxonomy of generalization
In Section 7, we introduce a taxonomy of generalization that extends beyond the coverage principle. We propose this taxonomy as an initial conceptual framework to disentangle different mechanisms of generalization that are often conflated in the literature. This categorization offers hypotheses for why neural networks succeed or fail at compositional tasks, connecting empirical observations to theoretical principles. We emphasize that the taxonomy is preliminary and that Sec. H.5 outlines its current limits, open questions, and future directions.
H.1 Type-I: Structure-based generalization
Type-I generalization occurs when a model succeeds by identifying and utilizing functionally equivalent components based on how primitive functions are composed. This form of generalization is precisely what we have formalized and studied via the coverage principle. Crucially, this kind of generalization remains strictly bounded by coverage as defined in Section 3.
In Type-I generalization, a model learns to recognize when different input fragments yield identical results in shared contexts, allowing it to generalize to new combinations of these fragments. However, without exposure to sufficient functional equivalence evidence during training, the model cannot reliably generalize to novel compositions.
Type-I generalization effectively describes what a pattern matching learner can achieve through the functional equivalence observed by input-output relationships, without requiring abstract variable binding mechanisms. This limitation aligns with the classic argument by Fodor and Pylyshyn [9] and Marcus [37] that connectionist networks lacking explicit variable binding capabilities [^13] We expand on the variable binding concept defined in Sec. 2 beyond flexibly associating variables with values, and we distinguish between self-replicative expansion (e.g., 2-hop $\to$ 3-hop via Type-III) and complex compositional binding (e.g., substituting noun phrases for nouns), where the latter may require explicit variable binding mechanisms beyond parameter sharing. cannot exhibit fully systematic generalization. We introduce this perspective as a conceptual framing of the coverage principle: it describes the ceiling of pattern-based generalization in neural networks that rely on pattern matching rather than symbolic variable binding.
H.2 Type-II: Function property-based generalization
Type-II generalization occurs when a model exploits specific properties intrinsic to individual primitive functions, enabling generalization beyond patterns observed during training. Unlike Type-I generalization which requires explicit evidence of functional equivalence, Type-II generalization leverages invariant properties that hold for all inputs of the primitive, making the function’s output fully determined by those invariances even in unobserved contexts.
We identify several sub-families of property-based generalization:
  • Algebraic invariances: commutativity, associativity, and similar properties
  • Group-theoretic structure: as in Fourier basis representations for modular arithmetic [63]
  • Input irrelevance: attention sparsity that ignores distractors [81]
The reversal curse [16] exemplifies the layered nature of compositional challenges across multiple generalization types. The coverage principle (Type-I) explains the fundamental failure: training on “$A$ is $B$” provides no functional equivalence evidence for “$B$ is $A^{-1}$”. Property-based approaches like bidirectional training [83] partially succeed by exploiting relationship invertibility (Type-II), using architectural modifications to learn inverse mappings from the same training data. However, recent work [42] reveals deeper representational binding constraints: even though models can exploit inverse mapping properties in idealized conditions (abstract concept-level tasks), they struggle to maintain consistent concept representations when moving to realistic surface-level predictions, where entities switch roles between subject and object positions. This demonstrates how real compositional tasks involve multiple mechanism layers, where achieving one capability may be insufficient due to representational binding limitations.
Type-II generalization fundamentally differs from Type-I in that it can, in principle, transcend the limitations imposed by the coverage principle, allowing models to generalize correctly to inputs that lie outside the coverage of their training data by leveraging invariant properties of the underlying functions. Systematic characterization of pure Type-II extrapolation remains limited and is an important direction for future work.
H.3 Type-III: Shared-operator generalization
Type-III generalization emerges through reuse of identical primitive functions across different computational positions, e.g., when $f_{1}=f_{2}$ in a multi-hop reasoning task. Unlike Type-I (bounded by coverage) and Type-II (exploiting algebraic properties), Type-III generalization leverages the structural repetition of identical operations.
This mechanism is not unique to Transformers. Recurrent architectures [84] exemplify Type-III generalization through their recurrent weight sharing: the same transformation is applied at each time step, enabling the network to process sequences of varying lengths despite training on fixed-length examples. This architectural inductive bias directly encodes the assumption that the same operation should apply across temporal positions, allowing recurrent architectures to extrapolate to longer sequences [85].
In the Transformer context, Wang et al. [19] demonstrate that Universal Transformer layers [86] enable generalization beyond in-domain data in two-hop tasks, while [87] show improved compositional reasoning with weight-shared layers. These findings suggest parameter sharing as a key enabling mechanism, where a single learned implementation of a function can be applied across multiple contexts.
A prominent manifestation of Type-III generalization appears in depth extrapolation and length generalization tasks, where models generalize to deeper or longer compositional structures than seen during training [99, 100, 101, 102]. While length generalization tasks fundamentally involve applying the same computation to longer inputs (Type-III), successful generalization often requires additional mechanisms like specialized positional encodings that facilitate computation reuse across positions. The observed brittleness of current approaches [103] suggests that such inductive biases are necessary but may not be sufficient for robust systematicity.
In summary, Type-III generalization extends beyond the coverage-bound limitations of Type-I by exploiting identical operations, without requiring the algebraic property understanding characteristic of Type-II. Type-III therefore complements Types I and II, but may still operate without explicit variable binding.
H.4 Relationship to prior taxonomies of generalization
Prior taxonomies of generalization in neural networks have primarily categorized the phenomena of compositional generalization rather than the underlying mechanisms. Lake and Baroni [12] distinguish between “mix-and-match” generalization (for small training-test differences) and systematic compositional generalization (requiring abstract rule application). Other frameworks focus on empirical evaluations through benchmarks [23] or decomposing compositional abilities into constituent skills [38].
Our taxonomy differs by categorizing the mechanisms that enable generalization rather than categorizing based on the phenomena. By distinguishing between structure-based pattern matching (Type-I), algebraic property exploitation (Type-II), and shared-operator transfer (Type-III), our framework provides a lens for interpreting why models succeed or fail at certain compositional tasks. This mechanistic approach not only explains observed behaviors but also offers hypotheses about when Types II or III become necessary for certain generalization challenges.
H.5 Open questions and future directions
To operationalize this taxonomy, future work could develop diagnostics: coverage ablation tests for Type-I (systematically removing training examples to test coverage boundaries), algebraic identity probes for Type-II (testing whether models recognize properties like commutativity without explicit training), and parameter-sharing ablations for Type-III (comparing shared vs. independent weights across positions). Current examples in this paper are illustrative rather than definitive classifications; real tasks often involve combinations of mechanisms, and our taxonomy provides a scaffold for disentangling them rather than enforcing mutually exclusive categories.
Several fundamental questions remain. First, practical computations often involve mixed ingredients of generalization, spanning multiple types. When classification is ambiguous, we suggest identifying the dominant mix of types that contribute, as most realistic tasks involve combinations rather than pure instances of any single type. Second, identifying the true compositional structure underlying a task can be ambiguous, as Locatello et al. [52] demonstrate regarding multiple valid decompositions. Third, real-world datasets contain heterogeneous computational structures, and it remains unclear how neural networks accommodate this diversity. Fourth, none of the three mechanisms solves the variable binding problem, and devising diagnostics and architectures for binding remains crucial.
Future work could explore connections to complementary frameworks:
  • Minimum Description Length [73]: Type-II may compress regularities that lower description length. Similarly, Type-III can be viewed as compression through parameter reuse.
  • Formal language learning [104, 105]: Connections to grammatical inference and automata learning
  • Cognitive science [106]: Human compositional learning strategies that may inspire architectural innovations
This taxonomy is not intended as a complete classification but rather as a conceptual framework to organize our understanding of generalization mechanisms and stimulate further research into the fundamental capabilities and limitations of neural networks on compositional tasks. Ultimately, developing architectural and optimization strategies that enable neural networks to leverage all three types more robustly and ultimately addressing the variable binding problem represents an important frontier for addressing the longstanding challenge of systematicity in connectionist systems.
Appendix I Partial computation observation drives the alignment of functional equivalence representation and vocabulary space
In this section, we investigate how exposure to partial computations affects the interpretability of intermediate state representations through vocabulary space alignment. We compare two training conditions on a modified 2-Hop task with $|\mathcal{X}|=50$ and $N=10\text{k}$, after 40k epochs of training:
  1. 1.
    Standard Training: $f_{1}\neq f_{2}$, model only sees complete two-hop examples $(x_{1},x_{2},x_{3})\mapsto t$.
  2. 2.
    With Partial Computation: $f_{1}=f_{2}$, model additionally sees all possible partial computations $(x_{1},x_{2})\mapsto b$ where $b=f_{1}(x_{1},x_{2})$ (2,500 partial examples, not counted toward the $N=10\text{k}$ two-hop training data).
To assess interpretability, we measure the Mean Reciprocal Rank (MRR) of intermediate state representations when projected to vocabulary space using the unembedding matrix. Low MRR indicates that the model’s internal representation of intermediate state $b$ aligns with the corresponding vocabulary token.
Fig. 19 shows a striking contrast between the two conditions. Under standard training, the MRR score remains very high throughout training, indicating that intermediate representations are not aligned with vocabulary space despite the model successfully learning the compositional task. However, when partial computations are included, the MRR score becomes very high, demonstrating clear vocabulary alignment.
Figure 19: MRR scores for intermediate state representations projected to vocabulary space. Left: Standard training ($f_{1}\neq f_{2}$, no partial computation) shows very high MRR regardless of position and layer. Right: Training with partial computation ($f_{1}=f_{2}$, with partial examples) shows MRR of 0 in layers 3 to 8 at position $x_{2}$, indicating strong vocabulary alignment.
This experiment suggests that logit lens interpretability is orthogonal to functional equivalence representation formation. A model can develop functionally correct intermediate representations that enable compositional generalization while remaining completely uninterpretable through standard vocabulary projection techniques. Interpretability via logit lens requires explicit vocabulary anchoring through exposure to partial computations that map intermediate states to vocabulary tokens.
This finding has important implications for mechanistic interpretability research: the absence of interpretable representations through logit lens does not indicate the absence of structured internal computation. Furthermore, it suggests that interpretability techniques may need to account for how training data shapes the alignment between internal representations and vocabulary space, rather than assuming such alignment emerges naturally from task performance.